P(X=0)=0.13

P(X=1)=0.17

P(X=2)=0.30

P(X=3)=0.32

P(X=4)=0.08

P(X=4)=0.08 is wrong and , i need help

Isabel Briggs Myers was a pioneer in the study of personality types. The personality types are broadly defined according to four main preferences. Do married couples choose similar or different personality types in their mates? The following data give an indication.

Similarities and Differences in a Random Sample of 375 Married Couples
Number of Similar Preferences     Number of Married Couples
All four 32
Three 120
Two 112
One 63
None 48

Suppose that a married couple is selected at random.

(a)

Use the data to estimate the probability that they will have 0, 1, 2, 3, or 4 personality preferences in common. (For each answer, enter a number. Enter your answers to 2 decimal places.)

0 1 2 3 4
Answer #1

Number of similar

preferences

Number of married couples probability
0 48 (48/375) =0.13
1 63 (63/375)= 0.17
2 112 (112/375)= 0.30
3 120 (120/375)= 0.32
4 32 (32/375)= 0.09
Total 375